# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ Abstract
The quantum expectation value and the stationary noise spectral density for a Fabri-Pérot gravitational-wave detector with a DC readout scheme are discussed in detail only through the quantum electrodynamics of lasers and the Heisenberg equations of mirrors’ motion. We demonstrate that the initial conditions of the mirrors’ motion concentrate around the fundamental frequency of the pendulum and are not related to the frequency range of our interest. Although in the ideal case, there is a consensus that the shot noise for the laser in the high-frequency range in the signal-referred noise spectral density decreases if the injected laser power is increased, our obtained noise spectral density shows that the shot noise does not decrease even if the injected laser power is increased. This is due to the leakage of the classical radiation pressure forces from the classical carrier field to the output port, and the classical carrier field is used as the reference in the DC readout scheme. Since the classical radiation pressure forces affect the mirror motion as a classical constant force, this classical constant force changes the equilibrium point of the pendulum of mirrors’ motion. To recover the ideal case, we must consider adjusting the interferometer’s tuning point to place the mirrors in equilibrium points. We investigate the case where the equilibrium tuning is incomplete and show that the behavior of the above shot noise is due to this incompleteness. We also discuss the maximal deviation of the mirror displacements from the equilibrium point in the incomplete tuning to recover the near ideal case.
@@ I Introduction
A decade ago, the first detection of the gravitational-wave signal GW150914 from a black hole-black hole binary was achieved. This event marked the beginning of gravitational-wave astronomy and multi-messenger astronomy including gravitational-wave detection [1, 2]. We are now at the stage where we can directly measure gravitational waves, and we can conduct scientific research through these gravitational-wave events. Many events, mainly from black hole-black hole binaries, have already been detected. We can also expect that one of the future directions of gravitational-wave astronomy is the development as a “precise science.” This will involve detailed studies of source science from the astrophysical point of view, the tests of general relativity, and the developments of the global network of gravitational-wave detectors [2, 3, 4, 5]. In addition to the current network of ground-based detectors, as future ground-based gravitational-wave detectors, the projects of Einstein Telescope [6] and Cosmic Explorer [7] are in progress, aiming to achieve more sensitive detections.
@@
To advance the field of gravitational-wave detections as a more precise science, it is essential to improve detector sensitivity. We note that theoretical developments in detector science are also crucial for enhancing our basic understanding of these gravitational-wave detectors. The current gravitational-wave detectors are limited by the shot noise, which is one of the quantum noises in the interferometer. In interferometric gravitational detectors, it is considered that there are two kinds of quantum noises. One is the shot noise of the laser, which constrains the sensitivity in the high-frequency range of the gravitational-wave detectors. On the other hand, the quantum radiation pressure forces are considered as a constraint on the sensitivity in the low-frequency region. In the common understanding in the gravitational-wave detection community, if the injected power is increased, the shot noise is decreased, while the quantum radiation pressure forces are increased [8]. The envelop of this trade-off relation in the signal-referred noise spectral density is regarded as the “standard quantum limit.” It is also a common understanding in the gravitational-wave detection community that this standard quantum limit is also estimated from the Heisenberg uncertainty relation of the noncommutativity of the quantum positions and the momenta of test masses [9].
@@
Furthermore, in the current research on the gravitational-wave detection, there are some reports which state that the current LIGO gravitational-wave detectors already violate the above “standard quantum limit” [10, 11]. Even from a theoretical perspective, this presents an opportune moment to revisit and refine the theoretical arguments, aiming for greater accuracy than those presented in previous works in Refs. [8, 9].
@@
On the other hand, a mathematically rigorous quantum measurement theory in quantum mechanics has also been developed (Ref. [12] and the references therein), which introduces new error-disturbance relations, distinct from the Heisenberg uncertainty relation. These new error-disturbance relations have already been confirmed through experiments [14, 15, 16]. One of the motivations of this development was the detection of gravitational-waves [13]. However, the actual application of this mathematical theory to the gravitational-wave detectors requires its extension to the quantum field theories, because the quantum noise in gravitational-wave detectors is analyzed through the quantum field theories of lasers [8]. Moreover, in the quantum measurement theory, it is essential to specify the final measured operator in the quantum measurement process due to the von Neumann chain problem [17]. In interferometric gravitational-wave detectors, the directly measured quantum operator is identified within the “readout scheme” of the detectors. While current gravitational-wave detectors utilize a feedback control system, and the final measured data of gravitational-wave detectors consist of the electric currents from this feedback control, the readout scheme in gravitational-wave detectors is the optical system that determines the finally measured quantum operator in the optical fields detected at the photodetectors, which receive signals from the main interferometer. Therefore, research into this readout scheme is crucial for advancing mathematical quantum measurement theory and its application to gravitational-wave detections.
@@
Due to the issue of the von Neumann chain, this paper adopts the perspective that the quantum properties in the laser interferometers are preserved until the photons are detected. At the photo-detectors, the quantum information of photons is transformed into an electric current. The processes that occur within the photodetectors are complex and vary depending on the type of device used. As a result, we speculate that quantum decoherence takes place during this photodetection process, and the resulting electric current, generated from the statistical results of the detection, can be regarded as classical currents. If this perspective is incorrect and the quantum nature is retained even in the feedback electric current, we would need to explore the concept of a quantum feedback control system [18]. This consideration is beyond the current scope of this paper.
@@
Current gravitational-wave detectors utilize a “DC readout scheme,” which is explained in Sec. III of this paper. Our first step is to investigate the DC-readout scheme from a quantum theoretical perspective. We previously discussed the mathematically rigorous quantum description of the balanced homodyne detection as a readout scheme in gravitational-wave detectors [19], since the balanced homodyne detection is planned for installation in future gravitational-wave detectors. In this paper, we will consider the DC readout scheme as a readout scheme for gravitational-wave detectors. Our investigation in this paper is a natural extension of the quantum theoretical arguments of a balanced homodyne detection [19]. The key aspect of our argument lies in the specification of the quantum operator that is finally observed. We assume that the finally observed quantum operator in the photodetectors is Glauber’s photon number [20], as discussed in Sec. III. It is often stated that the finally observed quantum operator is the power operator of the laser. However, even when we utilize the laser’s power operator, the conclusions we arrive at for the DC readout scheme are consistent with those related to Glauber’s photon number. Generally speaking, if we employ a different quantum operator, such as the mode-by-mode number operator, we may draw different conclusions [21]. This variation depends on the photodetection device, as mentioned above. Therefore, the specification of our finally measured quantum operator is essential to our arguments.
@@
After discussing the general arguments related to the DC readout scheme, we will examine the input-output relation for a Fabri-Pérot gravitational-wave detector. A specific input-output relation is essential for a detailed analysis of the DC readout scheme. We aim to derive this input-output relation of a Fabri-Pérot gravitational-wave detector, starting without specifying the small motions of the mirrors. These small motions of mirrors encapsulate information about external forces, which include gravitational-wave signals and the radiation pressure forces from the laser. In this paper, we describe the small motions of mirrors using the Heisenberg equation in quantum mechanics for a forced harmonic oscillator. While previous literature has often considered the mirrors’ small motions as free except for the gravitational-wave signal and the radiation pressure forces, we have introduced the fundamental frequency $\omega_{p}$ of the vibration isolation system of the pendulum supporting the mirrors.
@@
We find that the classical radiation pressure force also affects the mirror, which is finite due to the introduced fundamental frequency $\omega_{p}$ of pendulums. This classical radiation pressure manifests as a classical carrier field that leaks to the output port of the interferometer. This leakage influences the noise estimation in the DC readout scheme. We discuss the effects of this leakage of classical radiation pressure forces in the stationary noise spectral densities. Notably, when we consider the effects of this leakage, there is a situation where the shot noise contribution to the signal-referred noise spectral density in the high-frequency range does not decrease, even when the incident laser power is increased. To avoid the effects of the leakage of the classical radiation pressure forces, we have to adjust the tuning point of the Fabry-Pérot interferometer very carefully. We also estimate the maximum deviation from the ideal mirror position which realizes the near ideal case.
@@
Figure 1: The setup for the Fabry-Pérot interferometer includes notations for photoelectric quadratures. We define the classical distances $L$, $l_{x}$, and $l_{y}$, along with the small quantum displacements represented by $\hat{X}_{XEM}$, $\hat{X}_{XITM}$, $\hat{X}_{YEM}$, $\hat{X}_{YITM}$, $\hat{X}_{x}$, and $\hat{X}_{y}$ for the end mirrors (EMs) and the intermediate mirrors (ITMs). Additionally, the quadratures for the laser are denoted by $\hat{d}$ and $\hat{a}$ as the input quadratures to the interferometer, while $\hat{b}$ represents the output quadrature from the interferometer. The quadratures $\hat{c}_{x}$ and $\hat{c}_{y}$ are for the laser, which is separated by the beam splitter (BS), and $\hat{c}_{x}^{\prime}$ and $\hat{c}_{y}^{\prime}$ denote the quadratures that returned from the $x$- and $y$-cavities to the BS. The notation of the quadratures of the laser between the ITMs and EMS is illustrated in Fig. 2.
@@
We need to emphasize that we do not discuss the power recycling, the signal recycling, or the squeezed state input techniques [8, 9, 22, 23, 24] in this paper. Our focus is solely on re-evaluating a simple Fabri-Pérot gravitational-wave detector. Therefore, within this paper, we cannot discuss the recent findings regarding the “violation of the standard quantum limit” in LIGO gravitational-wave detectors [10, 11]. However, we expect that the arguments presented in this paper can be extended to include techniques such as power recycling, signal recycling, or squeezed state input techniques. In this regard, we are confident that the arguments in this paper are meaningful.
@@
The organization of this paper is as follows. In Sec. II, we summarize the basic notation that is used in this paper. In Sec. III, we develop the general arguments of the DC readout scheme. In Sec. IV, we derive the input-output relation without any specification of the small mirror displacements. In Sec. V, we consider the Heisenberg equations of motion to determine the small mirror displacements and their solution. In Sec. VI, we derive the final input-output relation by the specifications of the small mirror displacements through the Heisenberg equation of motion. We also discuss the comparison with the noise spectral densities in Ref. [8] also given in this section. In Sec. VII, we discuss the expectation value and the stationary noise spectral density in the DC readout from the results in Secs. III and VI. In this section, the resulting noise spectral density shows that the shot noise in the high-frequency region does not decrease even if the injected laser power is increased. In Sec. VIII, we consider changing the tuning point of the interferometer to recover the ideal noise spectral density in Ref. [8].
@@ II.1 Electric field notation
where $\hat{a}(\omega)$ is the photon annihilation operator associated with the electric field $\hat{E}_{a}(t-z)$, which satisfies the commutation relation
@@
$\cal A$} is the cross-sectional area of the optical beam. To discuss the input-output relation of the interferometer, based on one-photon formulation, it is convenient to introduce the operator $\hat{A}(\omega)$ defined by
@@
We note that the function $\Delta_{a}(t-t^{\prime})$ has an ultraviolet divergence. However, in the actual measurements of the time sequence of the variables, the time in a measurement is discrete with a finite time bin. This time bin gives the maximum frequency $\omega_{\max}$, which becomes the natural ultraviolet cut-off of the frequency in the obtained data. Incidentally, in the actual measurements of the time sequence of the variables, the whole measurement time is also finite. It gives the minimum frequency $\omega_{\min}$ which corresponds to a natural infrared cut-off in frequency. Therefore, we may regard that the integration range over $\omega$ in the definition of the function $\Delta_{a}(t-t^{\prime})$ in Eq. (13) is $\omega_{\min},\omega_{\max}]$ instead of $0,+\infty]$. For this reason, throughout this paper, we do not regard the divergence in the definition of the function $\Delta_{a}(t-t^{\prime})$ as a serious one. We apply similar arguments when we evaluate the averaged laser power $I_{0}$ in Sec. IV.5.
To discuss the quantum properties of the laser, we have to specify the quantum state of the electric field, which is expressed by Eqs. (1)–(3). One of the quantum states of the electric fields which considered within this paper is the vacuum state $|0\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
$\hat{a}(\omega)|0\rangle:=0\quad{}^{\forall}\omega>0.$ (14)
@@
$|\alpha\rangle_{a}$ $=$ $D_{a}[\alpha]|0\rangle_{a},$ (16)
$D_{a}[\alpha]$ $=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)\hat{a}^{\dagger}(\omega)-\alpha^{*}(\omega)\hat{a}(\omega)\right)\right].$
Here, we note that the subscription “aa” in the state $|0\rangle_{a}$ and $|\alpha\rangle_{a}$ indicate that these states are associated with the electric field operator $\hat{E}_{a}(t)$, with the quadrature $\hat{a}(\omega)$.

### v2

# Theoretical Detailed Analyses for DC readout and a Fabry-Pérot gravitational-wave detector

@@ Abstract
The quantum expectation value and the stationary noise spectral density for a Fabry-Pérot gravitational-wave detector with a DC readout scheme are discussed in detail only through the quantum electrodynamics of lasers and the Heisenberg equations of mirrors’ motion. We demonstrate that the initial conditions of the mirrors’ motion concentrate around the fundamental frequency of the pendulum and are not related to the frequency range of our interest. Although, in the ideal case, there is consensus that the shot-noise contribution from the laser to the high-frequency range of the signal-referred noise spectral density decreases as the injected laser power increases, our derived noise spectral density shows that the shot-noise contribution does not decrease. This is due to leakage of classical radiation pressure forces from the carrier field to the output port, and the carrier field is used as the reference in the DC readout scheme. Since classical radiation pressure acts as a constant force, it shifts the pendulum’s equilibrium point of the mirrors’ motion. To recover the ideal case, we must consider adjusting the interferometer’s tuning point to place the mirrors at their equilibrium positions. We investigate the case where the equilibrium tuning is incomplete and show that the behavior of the above shot noise is due to this incompleteness. We also discuss the maximum deviation of the mirror displacements from the equilibrium point during incomplete tuning to recover a near-ideal case.
@@ I Introduction
A decade ago, the first detection of the gravitational-wave signal GW150914 from a black hole-black hole binary was achieved. This event was the beginning of gravitational-wave astronomy and multi-messenger astronomy including gravitational-wave detection [1, 2]. We are now at a stage where we can directly measure gravitational waves and conduct scientific research using these events. Many events, mainly from black hole-black hole binaries, have already been detected. We can also expect that one of the future directions of gravitational-wave astronomy is to develop as a “precise science”, as we can easily extrapolate from the current achievements [3]. These involve detailed studies of source science from the astrophysical point of view, the tests of general relativity, and the developments of the global network of gravitational-wave detectors [2, 4, 5, 6]. In addition to the current network of ground-based detectors, as future ground-based gravitational-wave detectors, the projects of Einstein Telescope [7] and Cosmic Explorer [8] are in progress, aiming to achieve better sensitivity.
@@
To advance the field of gravitational-wave detections as a more precise science, it is essential to improve detector sensitivity. We note that theoretical developments in detector science are also crucial for enhancing our basic understanding of these gravitational-wave detectors. In the high-frequency part of the noise spectral density, current gravitational-wave detectors are limited by shot noise, one of the quantum noises in the interferometer. In interferometric gravitational detectors, it is considered that there are two kinds of quantum noises, although many plots in the gravitational-wave community do not distinguish these two quantum noises, e.g. see Ref. [9]. One is the shot noise of the laser, which limits the sensitivity of gravitational-wave detectors at high frequencies. On the other hand, quantum radiation pressure forces are considered a constraint on sensitivity in the low-frequency region. Note that, in the current experimental situation, this quantum radiation-pressure noise remains subdominant in this low-frequency region. In the common understanding in the gravitational-wave detection community, if the injected power is increased, the shot-noise contribution to the signal-referred noise spectral density is decreased, while the contribution of the quantum radiation pressure forces is increased [10]. The envelop of this trade-off relation in the signal-referred noise spectral density is regarded as the “standard quantum limit.” It is also a common understanding in the gravitational-wave detection community that this standard quantum limit is estimated from the Heisenberg uncertainty relation arising from the noncommutativity of the quantum positions and momenta of free test masses [11].
@@
Furthermore, in the current research on gravitational-wave detection, there are some reports which state that the current LIGO gravitational-wave detectors already surpassed the above “standard quantum limit” [12, 13]. Even from a theoretical perspective, this presents an opportune moment to revisit and refine the theoretical arguments, aiming for greater accuracy than those presented in previous works in Refs. [10, 11].
@@
On the other hand, a mathematically rigorous quantum measurement theory in quantum mechanics has also been developed (Ref. [14] and the references therein), introducing new error-disturbance relations distinct from the Heisenberg uncertainty relation. These new error-disturbance relations have already been confirmed through experiments [16, 17, 18]. One of the motivations of this development was the detection of gravitational waves [15]. However, the actual application of this mathematical theory to the gravitational-wave detectors requires its extension to the quantum field theories, because the quantum noise in gravitational-wave detectors is analyzed through the quantum field theories of lasers [10]. Moreover, in the quantum measurement theory, it is essential to specify the final measured operator in the quantum measurement process due to the von Neumann chain problem [19]. In interferometric gravitational-wave detectors, the directly measured quantum operator is identified within the detectors’ “readout scheme.” While current gravitational-wave detectors utilize a feedback control system, and the final measured data of gravitational-wave detectors consist of the electric currents from this feedback control, the readout scheme in gravitational-wave detectors is the optical system that determines the final measured quantum operator in the optical fields detected at the photodetectors, which receive signals from the main interferometer. Therefore, research into this readout scheme is crucial for advancing mathematical quantum measurement theory and its application to gravitational-wave detections.
@@
Due to the von Neumann chain issue, this paper adopts the perspective that the quantum properties of the photons in the laser interferometers are preserved until they are detected. At the photodetectors, the quantum information of photons is transformed into an electric current. The processes that occur within photodetectors are complex and vary depending on the device type. As a result, we speculate that quantum decoherence occurs during this photodetection process, and the resulting electric current, derived from the statistical results of the detection, can be regarded as a classical current. If this perspective is incorrect and the quantum nature is retained even in the feedback electric current, we would need to explore the concept of a quantum feedback control system [20]. This consideration is beyond the current scope of this paper.
@@
Current gravitational-wave detectors utilize a “DC readout scheme,” which is explained in Sec. III of this paper. Our first step is to investigate the DC-readout scheme from a quantum theoretical perspective. We previously discussed the mathematically rigorous quantum description of the balanced homodyne detection as a readout scheme in gravitational-wave detectors [21], since the balanced homodyne detection is planned for installation in future gravitational-wave detectors [22]. In this paper, we will consider the DC readout scheme for gravitational-wave detectors. Our investigation in this paper is a natural extension of the quantum theoretical arguments of a balanced homodyne detection [21]. The key aspect of our argument lies in the specification of the quantum operator that is finally observed. We assume that the final observed quantum operator in the photodetectors is Glauber’s photon number [23], as discussed in Sec. III. For the monochromatic case, we consider the number operator $\hat{a}^{\dagger}(\omega)\hat{a}(\omega)$ for the quantum field associated with the annihilation operator $\hat{a}(\omega)$. Glauber’s photon number is a natural extension of this single-mode number operator to the multi-mode electric field. It is often stated that the final observed quantum operator is the power operator of the laser. However, even when we utilize the laser’s power operator, the conclusions we arrive at for the DC readout scheme are consistent with those related to Glauber’s photon number. Generally speaking, if we employ a different quantum operator, such as the mode-by-mode number operator, we may draw different conclusions [24]. This variation depends on the photodetection device, as mentioned above. Therefore, the specification of our final measured quantum operator is essential to our arguments.
@@
After discussing the general arguments related to the DC readout scheme, we will examine the input-output relation for a Fabry-Pérot gravitational-wave detector. A specific input-output relation is essential for a detailed analysis of the DC readout scheme [25]. We aim to derive this input-output relation of a Fabry-Pérot gravitational-wave detector, starting without specifying the small motions of the mirrors. These small mirror motions encapsulate information about external forces, including gravitational-wave signals and radiation pressure from the laser. In this paper, we describe the small motions of mirrors using the Heisenberg equation in quantum mechanics for a forced harmonic oscillator [26]. While the previous literature [10, 11] has often considered the mirrors’ small motions as free except for the gravitational-wave signal and the radiation pressure forces, we have introduced the fundamental frequency $\omega_{p}$ of the vibration isolation system of the pendulum supporting the mirrors.
@@
We find that the classical radiation pressure force also acts on the mirror, which is finite due to the introduced fundamental frequency $\omega_{p}$ of the pendulums. This classical radiation pressure manifests as a classical carrier field that leaks to the output port of the interferometer. This leakage influences the noise estimation in the DC readout scheme. We discuss the effects of this leakage of classical radiation pressure forces in the stationary noise spectral densities. Notably, when we consider the effects of this leakage, there is a situation where the shot-noise contribution to the signal-referred noise spectral density in the high-frequency range does not decrease, even when the incident laser power is increased. To avoid the effects of the leakage of the classical radiation pressure forces, we have to adjust the tuning point of the Fabry-Pérot interferometer very carefully. We also estimate the maximum deviation from the ideal mirror position, which realizes the near ideal case.
@@
Figure 1: The setup for the Fabry-Pérot interferometer includes notations for photoelectric quadratures. We define the classical distances $L$, $l_{x}$, and $l_{y}$, along with the small quantum displacements represented by $\hat{X}_{XEM}$, $\hat{X}_{XITM}$, $\hat{X}_{YEM}$, $\hat{X}_{YITM}$, $\hat{X}_{x}$, and $\hat{X}_{y}$ for the end test masses (EMs) and the input test masses (ITMs). Additionally, the quadratures for the laser are denoted by $\hat{d}$ and $\hat{a}$ as the input quadratures to the interferometer, while $\hat{b}$ represents the output quadrature from the interferometer. The quadratures $\hat{c}_{x}$ and $\hat{c}_{y}$ are for the laser, which is separated by the beam splitter (BS), and $\hat{c}_{x}^{\prime}$ and $\hat{c}_{y}^{\prime}$ denote the quadratures that returned from the $x$- and $y$-cavities to the BS. The notation of the quadratures of the laser between the ITMs and EMs is illustrated in Fig. 2.
@@
We need to emphasize that we do not discuss the power recycling, the signal recycling, or the squeezed state input techniques [10, 11, 27, 28, 29] in this paper. Our focus is solely on re-evaluating a simple Fabry-Pérot gravitational-wave detector. Therefore, within this paper, we cannot discuss the recent findings regarding the “violation of the standard quantum limit” in LIGO gravitational-wave detectors [12, 13]. However, the main purpose of this paper is to point out two facts. One is the fact that the quantum uncertainty due to the motion of test masses does not affect the noise spectral density. The other is the fact the additional classical-carrier-field leakage due to the imperfect tuning of the interferometer at the equilibrium point of the pendulum affect the noise spectral density as an imperfection of the gravitational-wave detectors. The realization of the currently operating gravitational-wave detectors is beyond the current scope of this paper. Furthermore, we expect that the arguments presented in this paper can be extended to include techniques such as power recycling, signal recycling, and squeezed-state input. In this regard, we are confident that the arguments in this paper are meaningful.
@@
The organization of this paper is as follows. In Sec. II, we summarize the basic notation that is used in this paper. In Sec. III, we develop the general arguments of the DC readout scheme. In Sec. IV, we derive the input-output relation without any specification of the small mirror displacements. In Sec. V, we consider the Heisenberg equations of motion to determine the small mirror displacements and their solution. In Sec. VI, we derive the final input-output relation by the specifications of the small mirror displacements through the Heisenberg equation of motion. We also discuss the comparison with the noise spectral densities in Ref. [10] also given in this section. In Sec. VII, we discuss the expectation value and the stationary noise spectral density in the DC readout from the results in Secs. III and VI. In this section, the resulting noise spectral density shows that the shot-noise contribution to the high-frequency region of the signal-referred noise spectral density does not decrease even as the injected laser power is increases. In Sec. VIII, we consider changing the tuning point of the interferometer to recover the ideal noise spectral density in Ref. [10].
@@ II.1 Electric field notation
where $\hat{a}(\omega)$ is the photon annihilation operator associated with the electric field $\hat{E}_{a}(t-z)$, which satisfies the commutation relations
@@
$\cal A$} is the cross-sectional area of the optical beam. To discuss the input-output relation of the interferometer, based on one-photon formulation. Although the two-photon formulation [35, 36] is widely used in the gravitational-wave community, this formulation includes some approximations. Since we also want to examine the validity of the two-photon formulation itself in this paper, we derive the input-output relation through the one-photon formulation starting from the point without any approximations. In the one-photon formulation, it is convenient to introduce the operator $\hat{A}(\omega)$ defined by
@@
We note that the function $\Delta_{a}(t-t^{\prime})$ has an ultraviolet divergence in its integration when $\omega\rightarrow\infty$. However, in the actual measurements of the time sequence of the variables, the time in a measurement is discrete with a finite time bin. This time bin gives the maximum frequency $\omega_{\max}$, which becomes the natural ultraviolet cut-off of the frequency in the obtained data. If the upper frequency of the signal of our interest is 20kHz, $\omega_{\max}$ must be larger than $2\pi\times 20\times 10^{3}$Hz. Incidentally, in the actual measurements of the time sequence of the variables, the whole measurement time is also finite. It gives the minimum frequency $\omega_{\min}$ which corresponds to a natural infrared cut-off in frequency. If the lower frequency of the signal of our interest is 20Hz, $\omega_{\min}$ must be smaller than $2\pi\times 20$Hz. Therefore, we may regard that the integration range over $\omega$ in the definition of the function $\Delta_{a}(t-t^{\prime})$ in Eq. (13) is $\omega_{\min},\omega_{\max}]$ instead of $0,+\infty]$. For this reason, throughout this paper, we do not regard the divergence in the definition of the function $\Delta_{a}(t-t^{\prime})$ as a serious one. We apply similar arguments when we evaluate the averaged laser power $I_{0}$ in Sec. IV.5.
To discuss the quantum properties of the laser, we have to specify the quantum state of the electric field, which is expressed by Eqs. (1)–(3). One of the quantum states of the electric fields which is considered within this paper is the vacuum state $|0\rangle_{a}$ associated with the quadrature $\hat{a}(\omega)$, which is defined by
$\hat{a}(\omega)|0\rangle_{a}:=0\quad{}^{\forall}\omega>0.$ (14)
@@
$|\alpha\rangle_{a}$ $=$ ${{\mathfrak{D}}}_{a}[\alpha]|0\rangle_{a},$ (16)
${{\mathfrak{D}}}_{a}[\alpha]$ $=$ $\exp\left[\int\frac{d\omega}{2\pi}\left(\alpha(\omega)\hat{a}^{\dagger}(\omega)-\alpha^{*}(\omega)\hat{a}(\omega)\right)\right].$
Here, we note that the subscription “aa” in the state $|0\rangle_{a}$ and $|\alpha\rangle_{a}$ indicate that these states are associated with the electric field operator $\hat{E}_{a}(t)$ with the quadrature $\hat{a}(\omega)$.